• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 17:13:38 on July30,2003

    Entry number 108220

    keyword=L-arm DAQ working

    Jack replaced the broken CAMAC supply and the Left Arm DAQ works again.
    I downloaded the trigger, did some test runs including 500 Hz rate (pulser), and
    have started a cosmics run 23001.

    During Monday night's troubles a high deadtime was observed - it was because the
    gates passed through this defective crate. Absence of gate leads to long (3 msec)
    deadtime. In other words, ALL the symptoms were due to this crate.

    While checking today I noticed that the helicity on the redundant scaler did not
    agree with the primary helicity gated scaler. This was because of a flakey cable
    which probably was bumped during the effort to replace a scaler. (A model 1151
    scaler had been replaced by R. Feuerbach on Tuesday because of Bus Errors.)
    Now the redundant scalers agree again. I think all is ok.